OpenStudy (gabby1214):

solve the following equations 1. -5=x+3 2. w-2=-3 3. 10.5=5z 4. -8=z/3 5. -7=11/17z

pooja195 (pooja195):

1. -5=x+3 We are trying to get x alone correct? simply subtract 3 from both sides x=?

pooja195 (pooja195):

2. w-2=-3 Solving for w add 2 to both sides w=?

OpenStudy (gabby1214):

1. is -2? and 2. is -5??

pooja195 (pooja195):

1) correct 2) -3+2=?

OpenStudy (gabby1214):

isn't it 5 though

pooja195 (pooja195):

No we are adding to a negative number

OpenStudy (gabby1214):

so -5? I'm sorry I'm confused

pooja195 (pooja195):



pooja195 (pooja195):

look at -3 move 2 spaces to the right

OpenStudy (gabby1214):

Ohhhh okay I get that now thank you!

pooja195 (pooja195):

Wanna finish the rest?

OpenStudy (gabby1214):

yes please

pooja195 (pooja195):

ok :) 3. 10.5=5z here just divide both sides by 5 5/5=1 so just z 10.5/5=?

OpenStudy (gabby1214):

0.5?

pooja195 (pooja195):

Try again remeber that we are dividing

OpenStudy (gabby1214):

15.5? or just 5?

pooja195 (pooja195):

10.5 divided by 5

OpenStudy (gabby1214):

oh ok then its 2.1

pooja195 (pooja195):

Correct 4. -8=z/3 Multiply both sides by 3 -8 x 3=z

pooja195 (pooja195):

What is 8 times -3?

OpenStudy (gabby1214):

-24?

pooja195 (pooja195):

Good 5. -7=11/17z multiply both sides by 17/11 17/11 x -7 =?

OpenStudy (gabby1214):

77/17?

pooja195 (pooja195):

Nope try again.

OpenStudy (gabby1214):

the other way around? 17/77?

pooja195 (pooja195):

Use a calculator if needed

OpenStudy (gabby1214):

11/119?

pooja195 (pooja195):

Good =)

pooja195 (pooja195):

That's all it was

OpenStudy (gabby1214):

thank you so much!

pooja195 (pooja195):

yw ^_^